One canned juice drink is 30% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 25 liters that is 7% of orange juice.
Answer by stanbon(75887)

One canned juice drink is 30% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 25 liters that is 7% of orange juice.
----
Equation:
orange + orange = orange
0.30x + 0.05(25-x) = 0.07*25
--------
30x + 5*25 - 5x = 7*25
------
25x = 2*25
----
x = 2 liters (amt. of 30% juice needed)
25-x = 23 liters (amt. of 5% juice needed)
---------
Cheers,
Stan H.
------------